I am trying to work out a problem and I am not sure how to work it. The problem was given to my class out of the book for extra credit.
The problem is: 2x2-x-21.
It is 2xsquared subtract x subtract 21
Thank You
Brittany Miller

Answer by checkley71(8403)

2X~2-X-21=0 OR (2X-7)(X+3)=0 OR 2X-7=0 OR 2X=7 OR X=7/2 AND X+3=0 OR X=-3